OpenStudy (anonymous):

Solve on the interval [0,2Pi): A. Pi/3 , 2pi/3 B. Pi/6, 5pi/6 C. Pi/3, 5pi/3 D. 2pi/3, 4pi/3

OpenStudy (anonymous):

Equation is 1 +sin theta = (sqrt(3) + 2)/2

OpenStudy (anonymous):

@CliffSedge

OpenStudy (anonymous):

start with \[\sin(\theta)=\frac{\sqrt{3}}{2}\] and then it should be straight forward

OpenStudy (anonymous):

there is a plus 2 on top though?

OpenStudy (anonymous):

\[\frac{\sqrt{3}+2}{2}=\frac{\sqrt{3}}{2}+\frac{2}{2}\]

OpenStudy (anonymous):

er im getting confused

OpenStudy (anonymous):

did you understand my second post?

OpenStudy (anonymous):

why did you seperate it?

OpenStudy (anonymous):

so show you that the number on the right was identical to \[\frac{\sqrt{3}}{2}+1\] so that when you subtract 1 from both sides as a first step you get \[\sin(\theta)=\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

there isn't a plus one on that side...

OpenStudy (anonymous):

\[a+b=c+a\]\[a-a+b=c+a-a\]\[b=c\]It's just one of those rules

OpenStudy (anonymous):

wah? what does that have to do with it..

OpenStudy (anonymous):

This is your equation:\[1+\sin \theta=\frac{ \sqrt 3 }{ 2 }+1\]Treat one as a in my above response. That's what I meant--since you are adding the same number to both sides, you can subtract that same number from both sides to make the equation into simply sine of theta equals one half the square root of three

OpenStudy (anonymous):

there is no plus one on the right .... its just a divide 2...

OpenStudy (anonymous):

Did you see where satellite showed you how to separate the fraction giving a +2/2 at the end?

OpenStudy (anonymous):

no .i guess my pc being stupid? i dunno

OpenStudy (anonymous):

wait.... the +2 is seperate soyou can just cancel it out?

OpenStudy (anonymous):

Let's start over. \[\large Given \space \space 1 +\sin \theta = \frac{\sqrt{3} + 2}{2}\] \[\large 1 +\sin \theta = \frac{\sqrt{3}}{2} + \frac{2}{2}\] \[\large \rightarrow 1 +\sin \theta = \frac{\sqrt{3}}{2} + 1 \] \[\large \rightarrow \sin \theta = \frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

Do you understand all the steps I took to simplify that equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

What angle has a sine of √(3)/2?

OpenStudy (anonymous):

i can't figure out how to solve for the angle

OpenStudy (anonymous):

Consider some common right triangles.

OpenStudy (anonymous):

3 /4 /5 ?

OpenStudy (anonymous):

Ok, what others do you know?

OpenStudy (anonymous):

Sine is opposite-over-adjacent. 

OpenStudy (anonymous):

eh...idon't know any others tbh ...thats the only one i can think of right off the top of my head

OpenStudy (anonymous):

oh okay so thats a 30?

OpenStudy (anonymous):

i know the angle thing...

OpenStudy (anonymous):



OpenStudy (anonymous):

On a calculator, you use the inverse-sine function \[\sin^{-1}\left(\frac{\sqrt{3}}{2}\right) = 60º \]

OpenStudy (anonymous):

Now all you need to do is find the other angle that also has that sine on the interval you were given (and convert the angles to radians).

OpenStudy (anonymous):

interval ?

OpenStudy (anonymous):

what i mean is sin on an interval..its range of numbers?

OpenStudy (anonymous):

Trig functions are periodic (they repeat their values regularly). On the interval for Θ given [0,2π), there are two angles that have that same sine ratio.

OpenStudy (anonymous):

i can't graph it... how to graph it...

OpenStudy (anonymous):

Do you need to graph it? You can look at a unit circle to help find the other angles. Try http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

I understand up to sin theta = sqrt3/2

OpenStudy (anonymous):

@CliffSedge sorry to bother you with this

OpenStudy (anonymous):

What's up?

OpenStudy (anonymous):

I need to know how to find the angle and such..I am not sure i even understand the problem...._. can you perhaps do a step by step with me :/ @CliffSedge

OpenStudy (anonymous):

@hartnn help me please ._.

hartnn (hartnn):

'I understand up to sin theta = sqrt3/2' sure ?

OpenStudy (anonymous):

Did you look at the problem?

hartnn (hartnn):

yes, sin theta =sqrt 3 /2 is preety straight forward

hartnn (hartnn):

now for what theta is sin ratio = sqrt 3/2 ?

OpenStudy (anonymous):

yeah i guess i don't understand thequestion really ._.

hartnn (hartnn):

look for what angles is sin theta = root3/2

OpenStudy (anonymous):

30 degrees?

hartnn (hartnn):

sin 30 = 0.5

OpenStudy (anonymous):

and? ...

hartnn (hartnn):

so its not 30 lol! :P

OpenStudy (anonymous):

but there was a rt3/2 next to it?

hartnn (hartnn):

SEE THE Y-CO-ORDINATES , NOT X

OpenStudy (anonymous):

Omg...DUH! xD

OpenStudy (anonymous):

120 ? :3

hartnn (hartnn):

yes, and in 1st quadrant ?

OpenStudy (anonymous):

yeah that would be in the first quadrant

hartnn (hartnn):

in first quad its 60

OpenStudy (anonymous):

wait.... why we going to that?

hartnn (hartnn):

pi/3 and 2pi/3